OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <1, -2>, v = <-4, 8>

OpenStudy (e.mccormick):

I would start with the dot product. If it is 0, then they are orthogonal, and it is easy to do.

OpenStudy (anonymous):

(1)(-4) + (-2)(8) = -4 + (-16) = -20

OpenStudy (e.mccormick):

So that eliminates that right off.

OpenStudy (anonymous):

OKay so it can be parallel or neither

OpenStudy (e.mccormick):

Next, would be to look and see if they are parallel.

OpenStudy (e.mccormick):

If you can multiply by some scalar on one vector and get the other, they are parallel.

OpenStudy (e.mccormick):

So, can you find a c such that \(c\vec{u}=\vec{v}\)

OpenStudy (anonymous):

Idk what that means lol... what do i do?

OpenStudy (amistre64):

just stack the vectors, if they result in a common ratio they are paraleel

OpenStudy (e.mccormick):

Ah, do you know what scalar multiplication of a vector is? If not, I can show that real fast.

OpenStudy (amistre64):

u = <1, -2> v = <-4, 8> -1/4 = -2/8

OpenStudy (e.mccormick):

@amistre64 Ah, that is a nice test for it.

OpenStudy (anonymous):

So since they are a common ratio then that means they are parallel right?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

Thanks guys! how do i give a medal to both of you?

OpenStudy (amistre64):

i have enough medals :)

OpenStudy (anonymous):

Oh true lol

OpenStudy (e.mccormick):

That can also be seen with \(-4\vec{u}=\vec{v}\) but that method is much nicer!

OpenStudy (anonymous):

Thanks again guys

OpenStudy (e.mccormick):

np. Have fun!